OpenStudy (anonymous):

Solve using the multiplication principle. -18 = 19y please help

OpenStudy (chriss):

y=-18/19

OpenStudy (anonymous):

how did you get that answer

OpenStudy (chriss):

To isolate y, we have to divide 19y by 19. If we divide that side by 19, we have to do the same thing to the other side. -18=19y \[\frac{-18}{19}=\frac{19y}{19}\] \[\frac{-18}{19}=y\]

OpenStudy (anonymous):

thanks a lot

OpenStudy (chriss):

you're welcome